What come Next in the series 98,14,2,2/7
Answer by jim_thompson5910(35256)

The rule is that you multiply each term by 1/7 to get the next term. Multiply the last term by 1/7 to get (2/7)*(1/7) = (2*1)/(7*7) = 2/49